OpenStudy (kirbykirby):

sqrt(x) can be written as x^(1/2) We know (x^(a))^b = x^(ab) So, (x^2)^(1/2) = x^1 = x But, we know sqrt(x^2) = |x| What's this discrepancy about?

OpenStudy (kirbykirby):

\[\sqrt(x) =x ^{1/2} \] and \[(x^a)^b=x ^{ab}\]So, \[(x^{2})^{1/2}=x^1=x\] But we know that \[\sqrt{x^2}=\left| x \right|\]

ganeshie8 (ganeshie8):

whats the question

OpenStudy (kirbykirby):

Why does one method give x, and the other |x| ?

ganeshie8 (ganeshie8):

you mean, \((\sqrt{x} )^2 = x\) \((\sqrt{x^2} ) = |x|\)

ganeshie8 (ganeshie8):

why is it so, is that the question ?

OpenStudy (kirbykirby):

No, not exactly. Because \[\sqrt{x}=x^{1/2}\] so \[\sqrt{x^2}=(x^2)^{1/2}\]

OpenStudy (kirbykirby):

\[(\sqrt{x})^2=(x^{1/2})^2\]

ganeshie8 (ganeshie8):

wait

OpenStudy (kirbykirby):

Just the absolute bar is driving me nuts when we write sqrt(x^2) = |x|

ganeshie8 (ganeshie8):

whats domain of \((\sqrt{x})^2\) ?

ganeshie8 (ganeshie8):

and, whats domain of \(\sqrt{x^2}\) ?

ganeshie8 (ganeshie8):

if u understand why the domains are not same for both above, you wud see easily why absolute bars are coming for \(\sqrt{x^2}\)

OpenStudy (kirbykirby):

I guess x >=0 and x is real respectively ?

ganeshie8 (ganeshie8):

^ right

OpenStudy (kirbykirby):

well I kind of understand it when they use the square root sign. But I get puzzled when we say (x^a)^b= x^(ab), so (x^2)^(1/2) = x which disregards the absolute value.. o.o?

OpenStudy (kirbykirby):

I never saw a prof using absolute values when they wrote square roots as fractions

OpenStudy (zarkon):

assume \(x\ge 0\) then \((x^2)^{1/2}=x\) and \(((-x)^2)^{1/2}=(x^2)^{1/2}=x\) so for any \(x\) \[(x^2)^{1/2}=|x|\]

Parth (parthkohli):

Also notice,\[\rm \sqrt{(-x)^2} =\left( (-x)^2\right)^{1 \over 2} = (x^2)^{1\over 2} = x \]Now read that backwards. :)